NOTE: You are NOT required to write a program for this problem. We wish to determine whether at least one three-of-a-kind exists in a hand of N cards, where a three-of-a-kind is three of the same value. Assume the cards have values 1 – 13. For example, the hand with card values {11, 5, 11, 5, 5} does have at least one
three-of-a-kind. For each of the following algorithms (A and B), find the simplified total Big-O run time as a function of the number of cards (N). Explain your answers based on each of the bulleted items and how they contribute to the total run time.
Algorithm A:
• Create a “count” array of size N that stores, for each card in the hand, the number of cards with that card’s value. It is initialized to be filled with 1’s because we know there is at least 1 of each card.
• For each card in the hand:
o Compare its value to the value of every other card.
o Each time the two values match, increment the corresponding “count” element for this card.
• If at least one value in the final “count” array is 3 or greater, there is a three-of-a-kind because at least 3 cards match that given card’s value.
Example: With card values {11, 5, 11, 5, 5}, the values in the “count” array would become {2, 3, 2, 3, 3}, and there is at least one three-of-a-kind.
Algorithm B:
• Create a “count” array of size 13 that stores, for each possible card value, the number of cards in the hand that have that value (1 – 13). It is initialized to be filled with 0’s.
• For each card in the hand with value v, use v-1 as an index to increment the corresponding element in “count”. For example, if a card has value 13, increment “count” element #12.
• If at least one value in the final “count” array is 3 or greater, there is a three-of-a-kind because at least 3 cards have that given value.
Example: With card values {11, 5, 11, 5, 5}, the values in the “count” array would become {0, 0, 0, 0, 3, 0, 0, 0, 0, 0, 2, 0, 0}, and there is at least one three-of-a-kind.
Answer #1

Algorithm A: O(n^2) time + O(n) space
Algorithm B: O(n) time + O(1) space

Here is the step-wise break down.


Algorithm A:
• Create a “count” array of size N that stores, for each card in the hand, the number of cards with that card’s value. It is initialized to be filled with 1’s because we know there is at least 1 of each card. - takes O(n) space. O(n) time in initializing each value to 1
• For each card in the hand: - loop runs O(n) times
o Compare its value to the value of every other card. - O(n) comparisons made. Thus O(n)*O(n)
o Each time the two values match, increment the corresponding “count” element for this card. - O(1) time in incrementing counters
• If at least one value in the final “count” array is 3 or greater, there is a three-of-a-kind because at least 3 cards match that given card’s value. - O(n) time in searching for a value greater than or equal to 3.

Thus total time = O(n) + O(n)*[O(n)+O(1)] + O(n)
= O(n) + O(n^2) = O(n^2)


Algorithm B:
• Create a “count” array of size 13 that stores, for each possible card value, the number of cards in the hand that have that value (1 – 13). It is initialized to be filled with 0’s. - Constant time and space required for this. O(1)
• For each card - loop runs O(n) times - in the hand with value v, use v-1 as an index to increment the corresponding element in “count”. - Constant time lookup to increment counters. Thus O(n) * O(1)
• If at least one value in the final “count” array is 3 or greater, there is a three-of-a-kind because at least 3 cards have that given value. - O(n) time search to find a value at least 3

Thus total time = O(1) + O(n)*O(1) + O(n) = O(n)